OpenStudy (anonymous):

solve the equation for the interval [0,2pi) (secx)^2 - 2 = (tanx)^2

OpenStudy (raden):

i think there is somethink error with that equation ! please, recheck again ur problem ?

OpenStudy (anonymous):

\[\sec ^{2}x-2=\tan ^{2}x\]

OpenStudy (raden):

yea, i see.... but, do u know that (secx)^2 = (tanx)^2+1 what happend if we convert (secx)^2 to (tanx)^2+1 ?

OpenStudy (anonymous):

the problem is # 8 on the attachment I am going to put up

OpenStudy (raden):

well, from that equation : (secx)^2 - 2 = (tanx)^2 because (secx)^2 = (tanx)^2+1, so left hand side it can be (tanx)^2+1-2=(tanx)^ (tanx)^2-1=(tanx)^ (tanx)^2-(tanx)^2=1 0=1 so, no solution for x here...

OpenStudy (raden):

the answer is B

OpenStudy (anonymous):

could you look at #6 I got A for it. Is that correct?

OpenStudy (raden):

sorry i was answer other member question, wait i will check it

OpenStudy (raden):

A incorrect

OpenStudy (raden):

sorry, i was check all option.. no option are correct (if i no mistake calculate it)